OpenStudy (onepieceftw):

Prove: cos θ - cos θ•sin^2 θ = cos^3 θ. Can someone walk me through this?

OpenStudy (anonymous):

sin^2 +cos^2 =1 => sin^2=1-cos^2

OpenStudy (er.mohd.amir):

take co common and 1-sin^2 x=cos^2x

OpenStudy (anonymous):

cos - cos *(1-cos^2)=cos-cos +cos^3=cos^3

OpenStudy (onepieceftw):

Okay so sin^2 = 1-cos^2?

OpenStudy (anonymous):

yes

OpenStudy (onepieceftw):

Thank you so much!

OpenStudy (anonymous):

YW